NEW PERSPECTIVE ON PF_{n} (n=1–5) FROM THE RECOUPLED PAIR BONDING MODEL: A QUANTUM CHEMICAL STUDY

\underline{D. E. WOON} and T. H. DUNNING, JR., Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

Structures of the PF_{n} family (n=1–5) were characterized with high level RCCSD(T) coupled cluster theory calculations using triple and quadruple zeta quality correlation consistent basis sets. In addition to accounting for the well-known ground states of PF through PF_{5}, insight from the recoupled pair bonding model also led to locating a previously unknown 3B_{1} state of PF_{3}, which lies about 90 kcal/mol above PF_{3}(^{1}A_{1}) but is still bound with respect to PF_{2}(^{2}B_{1})+F(^{2}P) by about 40 kcal/mol. We also revisited the less-studied C_{3v} local minimum on the PF_{4} doublet surface and characterized the transition state for interconversion to the C_{2v} global minimum. The energetics suggest that both PF_{3}(^{3}B_{1}) and C_{3v} PF_{4}(^{2}A_{1}) are potentially observable in the laboratory. The trends in the bond dissociation energies and relative energy differences of the PF_{n} family are very consistent with predictions from the recoupled pair bonding model.